# Additive Models Explained: A Computational Complexity Approach

### v1

@@ 1 Introduction
Generalized additive models (GAMs) are a widely utilized family of models in ML (hastie2017generalized, ; nelder1972generalized, ), valued commonly for their interpretability (molnar2020interpretable, ; caruana2015intelligible, ; zhong2023exploring, ; bechler2024intelligible, ; changnode, ; liu2022fast, ; bordt2023shapley, ; enouen2025instashap, ; chen2020sparse, ; mctavish2024interpretable, ; mueller2024gamformer, ; enouen2022sparse, ; zhuang2021interpretable, ) due to their additive structure. Beyond being considered interpretable, the very assumption of “additive interpretability” is a foundational one in the explainable AI community. For instance, many explanation techniques are designed to approximate additive behaviors (ribeiro2016should, ), and metrics such as infidelity (yeh2019fidelity, ) assess explanation quality by measuring how well an additive approximation captures the effects of feature perturbations on the model’s output. Furthermore, interpretable training approaches often seek to promote near-additive behavior in models within specific domains (alvarez2018towards, ).
Since GAMs are commonly regarded as an interpretable class of ML models, it may seem natural to expect that different types of explanations for their predictions can be computed efficiently, without encountering drastic computational intractability (e.g., NP-hardness, #\#P-hardness, etc.). This question also relates to a growing body of research in the ML community focused on understanding the computational complexity involved in generating different types of explanations for various model classes (BaMoPeSu20, ; waldchen2021computational, ; bassanlocal, ). These efforts span both inherently complex models like neural networks (adolfi2024computational, ) and tree ensembles (audemard2022preferred, ), where explanations are expected to be computationally hard to compute (BaMoPeSu20, ; adolfi2024computational, ), as well as traditionally “interpretable” models such as decision trees (arenas2022computing, ) and monotonic classifiers (marques2020explaining, ), where computing explanations is expected to be more efficient (BaMoPeSu20, ; marques2023no, ; bassanlocal, ).
@@
    Component model type used within the GAM: Here, we investigate several popular component models that are typically used in GAMs and span the interpretability spectrum, including: (i) Smooth GAMs— i.e., the classical use of splines (hastie2017generalized, ); (ii) Neural Additive Models (NAMs)(agarwal2021neural, )— where each component is a neural network; and (iii) Explainable Boosting Machines (EBMs)(caruana2015intelligible, )— which use boosted tree ensembles as their components.
@@ 2 Preliminaries
Complexity Classes. This paper assumes familiarity with basic complexity classes such as polynomial-time (PTIME), and non-deterministic polynomial-time (NP, coNP). We also discuss the class $\Sigma_{2}^{P}$, which includes problems solvable in NP with access to a constant-time coNP oracle. Clearly, NP and coNP are contained in $\Sigma_{2}^{P}$, and it is also widely believed that NP, coNP $\subsetneq$ $\Sigma_{2}^{P}$ arora2009computational . Additionally, we cover #P, the class of functions that count the accepting paths of polynomial-time nondeterministic Turing machines. Although not directly comparable since they focus on counting rather than decision problems, it is widely believed that the #P class is strictly “harder” than $\Sigma_{2}^{P}$, which can be denoted as $\Sigma_{2}^{P}$ $\subsetneq$#P arora2009computational .
@@ 3.1 Dimension 1: The input domain
We study the following cases for the domain $\mathcal{X}_{i}$ of feature $i\in\{1,\ldots n\}$: (i) when the domain is enumerable discrete, meaning each value $\textbf{x}_{i}$ is selected from a given set $\mathcal{K}\subseteq\mathbb{Q}$ of a constant number $|\mathcal{K}|$ of possible assignments. Intuitively, these are scenarios where each input is limited to a predefined set of values, which is often relevant for cases such as tabular data or certain language tasks; (ii) the discrete setting, or the general discrete setting, where $\mathcal{X}_{i}$ represents any discrete input within a range defined by maximum and minimum values expressed in binary with $q$ bits. This intuitively allows a broader range of discrete assignments for each input, while still operating within a discrete domain. (iii) The continuous setting, where $\mathcal{X}_{i}\subseteq\mathbb{R}$, which ensures guarantees across the entire infinite spectrum of the specified domain . Both discrete and continuous domains are widely used across various ML tasks, including those in vision and language tasks (elkabetz2021continuous, ; hong2022bridging, ; cartuyvels2021discrete, ). We note that general discrete settings naturally encompass enumerable ones and are thus always at least as computationally challenging, whereas continuous settings do not enable a direct comparison to either and can make problems either harder or easier to solve (karmarkar1984new, ; vavasis1991nonlinear, ; arora2009computational, ).
@@ 3.2 Dimension 2: The component model type
We consider a diverse range of component models that can be integrated into a GAM, spanning various levels of interpretability and reflecting commonly used approaches in the literature: (i) The classic spline-based models (hastie2017generalized, ), where each component function is piecewise polynomial. In practice, these polynomials are typically restricted to a degree of at most three, yielding the widely used cubic splines (hastie2017generalized, ), which will be our primary focus (though our results can often be extended to higher-order splines — see Appendix B). To distinguish these GAMs from the non-smooth variants, we refer to GAMs with spline-based components as Smooth GAMs. (ii) Neural Additive Models (NAMs)(agarwal2021neural, ; bouchiat2023improving, ; xu2023sparse, ; jiao2024naisr, ; thielmann2024neural, ), in which each component is a neural network with ReLU activations. (iii) Explainable Boosting Machines (EBMs)(caruana2015intelligible, ; wei2022safety, ; nori2021accuracy, ), where each component function is a boosted tree ensemble. We present the complete formal definitions of all model types in Appendix B.
@@ 3.3 Dimension 3: The explanation type
To explore various dimensions of GAM interpretability, we examine a broad range of widely used explanation types drawn from the existing literature. Building upon prior computational complexity frameworks for evaluating explanations in ML models (BaMoPeSu20, ; arenas2021foundations, ; bassanlocal, ), we formalize each type of explanation as an explainability query. Such a query takes a model $f$ and x as inputs and aims to address specific questions, offering a meaningful interpretation of the prediction $f(\textbf{x})$.
Sufficient reason Feature Selection. We consider the common sufficiency criterion for feature selection, used in popular explainability methods (ribeiro2018anchors, ; carter2019made, ; ignatiev2019abduction, ). A sufficient reason is a subset of input features, $S\subseteq[n]$, such that fixing the features in $S$ to their values in $\textbf{x}\in\mathbb{F}$ guarantees the prediction remains $f(\textbf{x})$, regardless of the assignment to $\overline{S}$. We write $(\mathbf{x}_{S};\mathbf{z}_{\bar{S}})$ for an input where $S$ takes values from x and $\overline{S}$ from z. Formally, $S$ is a sufficient reason for $\langle f,\textbf{x}\rangle$ iff for all $\textbf{z}\in\mathbb{F}$: $f(\textbf{x}_{S};\textbf{z}_{\bar{S}})=f(\textbf{x})$.
A common assumption in the literature is that smaller sufficient reasons (those with smaller $|S|$) are more useful (ribeiro2018anchors, ; carter2019made, ; ignatiev2019abduction, ). This motivates the search for cardinally minimal sufficient reasons, also called minimum sufficient reasons, and leads to our first explainability query:
@@
To deepen our understanding of the complexity of sufficient reasons, we also examine two common related explainability queries that refine the MSR query (BaMoPeSu20, ; bassanlocal, ): (i) Check-Sufficient-Reason(CSR), which checks if a given subset $S$ is a sufficient reason; (ii) Count-Completions(CC), a generalized version of CSR that measures the fraction of completions preserving the prediction, capturing the probability of maintaining the classification. Full formalizations appear in Appendix C.
Contrastive Explanations. An alternative way to interpret models is by identifying subsets of features that, when changed, could alter the model’s prediction (dhurandhar2018explanations, ; guidotti2024counterfactual, ). We define s subset $S\subseteq[n]$ as contrastive if modifying its values can change the classification $f(\textbf{x})$, i.e., there exists $\textbf{z}\in\mathbb{F}$ such that $f(\textbf{x}_{\bar{S}};\textbf{z}_{S})\neq f(\textbf{x})$. As with sufficient reasons, smaller contrastive subsets are typically assumed to be more meaningful (dhurandhar2018explanations, ; guidotti2024counterfactual, ; ignatiev2020contrastive, ; mothilal2020explaining, ), motivating a focus on cardinally-minimal contrastive reasons:
@@
Shapley Values. In the additive attribution setting, each feature $i\in[n]$ is assigned an importance weight $\phi_{i}$. A common method for assigning these weights is the Shapley value attribution index (lundberg2017unified, ):
@@
We use the standard conditional expectation value function $v(S):=\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{z})\mid\textbf{z}_{S}=\textbf{x}_{S}]$ (sundararajan2020many, ; lundberg2017unified, ). Moreover, as is standard in both complexity analyses (arenas2023complexity, ; van2022tractability, ) and practical SHAP methods like KernelSHAP (lundberg2017unified, ), we assume feature independence. See Appendix C for a full formalization.
@@
Feature redundancy. For our final explanation form, we study the complexity of determining whether a feature $i$ is redundant with respect to a model $f$, following the notion of redundancy explored in prior work (arenas2021foundations, ; bassanlocal, ; huang2023feature, ; darwiche2020reasons, ). Unlike the previous local explanation forms, this one is global — it checks if $i$ is redundant for all $\textbf{x}\in\mathbb{F}$. Formally, $i$ is redundant if for all $\textbf{x},\textbf{z}\in\mathbb{F}$, we have $f(\textbf{x})=f(\textbf{x}_{[n]\setminus\{i\}};\textbf{z}_{\{i\}})$.
@@ 4 Main Complexity Results
NAMs, EBMs coNP-C NP-C coNP-C #\#P-C #\#P-C
@@
NAMs, EBMs coNP-C NP-C coNP-C #\#P-C #\#P-C
@@ 5 A Computational Interpretability Hierarchy for GAMs
Our results uncover a rich and varied landscape of complexity outcomes across different settings, which we analyze in this section. To present these complexity distinctions more elegantly, we adopt the notation from (BaMoPeSu20, ), indicating when one model is strictly more computationally interpretable (c-interpretable) than another — that is, when it belongs to a strictly lower complexity class.
@@ 5.2 The Significance of the Component Models on Complexity Depends on the Input Domain
We also note that the result shown in this subsection, together with our findings in Subsection 5.1 on input domain sensitivity, offers a complementary perspective to empirical work such as (chang2021interpretable, ). These studies demonstrate that both the choice of component model and the form of data representation (i.e., input representation) substantially influence the learned functional structure of GAMs. Since differing functional forms naturally correspond to different levels of computational complexity, this aligns with our observation that the complexity of generating explanations is highly sensitive to these modeling choices.
@@ 5.3 Complexity Distinctions Between Regression and Classification, but Only for SHAP
We note that while SHAP allows for a direct comparison between regression and classification tasks, adapting our other explanation types — originally defined for classification — to regression requires redefining them (as in (wu2024verix, ; izyaxuanjoigalma24, )) to assess whether values lie within a specified $\delta$-range. From a complexity standpoint, these behave similarly to classification cases and do not produce the same distinctions observed with SHAP. We elaborate on this in Appendix D.
@@
As noted in the introduction, the ML community often assumes that increasing a model’s additivity improves interpretability. For instance, NAMs (agarwal2021neural, ; radenovic2022neural, ) and EBMs (caruana2015intelligible, ) are generally seen as more interpretable than standard neural networks and boosted trees, respectively. It thus seems natural to expect that explanations for additive models (e.g., NAMs) are easier to compute than their non-additive counterparts. Our findings support this, showing a clear complexity gap between black-box models and their additive counterparts. However, importantly, this gap is not universal — it strongly depends on the explanation type and the input domain. Table 2 summarizes prior complexity results for neural networks and tree ensembles, alongside our new results for NAMs and EBMs.
@@
MSR PTIME $\mathbf{\Sigma^{P}_{2}}$-C coNP-C $\mathbf{\Sigma^{P}_{2}}$-C
@@ 6.1 Polynomial-Time Explanation Computations
Sufficient and contrastive reasons. We begin by analyzing the polynomial-time complexity of computing sufficient and contrastive reason queries (CSR, MSR, MCR). A key observation is that by identifying the minimal and maximal values each component $\beta_{i}\cdot f_{i}(\textbf{x}_{i})$ can take, we can rank features by their “importance”, measured by how much altering each feature deviates the prediction. A greedy algorithm, akin to those in (BaMoPeSu20, ; marques2020explaining, ), then selects features by importance until a minimal sufficient or contrastive reason is found. CSR is simpler, requiring only a sufficiency check for a given subset. Thus, the main computational challenge lies in efficiently computing the min/max values of each $f_{i}(\textbf{x}_{i})$. For enumerable discrete inputs, this can be done by enumeration; for smooth GAMs, these values can be computed analytically, even with general discrete or continuous inputs.
@@
SHAP explanations in regression settings. When computing SHAP, polynomial-time computation is feasible only for GAMs in regression tasks (classification intractability is discussed in Subsection 6.3). Leveraging the linearity axiom of Shapley values (lundberg2017unified, ), we reduce the computation to evaluating the computations of $\beta_{i}\cdot\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f_{i}(\textbf{x}_{i})]$. When the model is either discretely enumerable or supports efficient integration (e.g., in Smooth GAMs), these expectations can be computed in polynomial time.

### v2

@@ 1 Introduction
Generalized additive models (GAMs) are a widely utilized family of models in ML (Hastie, 2017; Nelder and Wedderburn, 1972), valued commonly for their interpretability (Molnar, 2020; Caruana et al., 2015; Zhong et al., 2023; Bechler-Speicher et al., 2024; Chang et al., 2022; Liu et al., 2022; Bordt and von Luxburg, 2023; Enouen and Liu, 2025; Chen et al., 2020; McTavish et al., 2024; Mueller et al., 2024; Enouen and Liu, 2022; Zhuang et al., 2021) due to their additive structure. Beyond being considered interpretable, the very assumption of “additive interpretability” is a foundational one in the explainable AI community. For instance, many explanation techniques are designed to approximate additive behaviors (Ribeiro et al., 2016), and metrics such as infidelity (Yeh et al., 2019) assess explanation quality by measuring how well an additive approximation captures the effects of feature perturbations on the model’s output. Furthermore, interpretable training approaches often seek to promote near-additive behavior in models within specific domains (Alvarez Melis and Jaakkola, 2018).
Since GAMs are commonly regarded as an interpretable class of ML models, it may seem natural to expect that different types of explanations for their predictions can be computed efficiently, without encountering drastic computational intractability (e.g., NP-hardness, #\#P-hardness, etc.). This question also relates to a growing body of research in the ML community focused on understanding the computational complexity involved in generating different types of explanations for various model classes (Barceló et al., 2020; Wäldchen et al., 2021; Bassan et al., 2024). These efforts span both inherently complex models like neural networks (Adolfi et al., 2025) and tree ensembles (Audemard et al., 2022), where explanations are expected to be computationally hard to compute (Barceló et al., 2020; Adolfi et al., 2025), as well as traditionally “interpretable” models such as decision trees (Arenas et al., 2022) and monotonic classifiers (Marques-Silva et al., 2020), where computing explanations is expected to be more efficient (Barceló et al., 2020; Marques-Silva and Ignatiev, 2023; Bassan et al., 2024).
@@
    Component model type used within the GAM: Here, we investigate several popular component models that are typically used in GAMs and span the interpretability spectrum, including: (i) Smooth GAMs— i.e., the classical use of splines (Hastie, 2017); (ii) Neural Additive Models (NAMs)(Agarwal et al., 2021)— where each component is a neural network; and (iii) Explainable Boosting Machines (EBMs)(Caruana et al., 2015)— which use boosted tree ensembles as their components.
@@ 2 Preliminaries
Complexity Classes. This paper assumes familiarity with basic complexity classes such as polynomial-time (PTIME), and non-deterministic polynomial-time (NP, coNP). We also discuss the class $\Sigma_{2}^{P}$, which includes problems solvable in NP with access to a coNP oracle and $\Delta^{P}_{2}$, which includes problems solvable in PTIME with access to an NP oracle. Clearly, NP and coNP are contained in $\Delta^{P}_{2}$ and $\Sigma_{2}^{P}$, and it is also widely believed that NP, coNP $\subsetneq\Delta^{P}_{2}\subsetneq$ $\Sigma_{2}^{P}$ Arora and Barak (2009). Additionally, we cover #P, the class of functions that count the accepting paths of polynomial-time nondeterministic Turing machines. Although not directly comparable since they focus on counting rather than decision problems, it is widely believed that the #P class is strictly “harder” than $\Sigma_{2}^{P}$, which can be denoted as $\Sigma_{2}^{P}$ $\subsetneq$#P Arora and Barak (2009).
@@ 3.1 Dimension 1: The input domain
We study the following cases for the domain $\mathcal{X}_{i}$ of feature $i\in\{1,\ldots n\}$: (i) when the domain is enumerable discrete, meaning each value $\textbf{x}_{i}$ is selected from a given set $\mathcal{K}\subseteq\mathbb{Q}$ of a constant number $|\mathcal{K}|$ of possible assignments. Intuitively, these are scenarios where each input is limited to a predefined set of values, which is often relevant for cases such as tabular data or certain language tasks; (ii) the discrete setting, or the general discrete setting, where $\mathcal{X}_{i}$ represents any discrete input within a range defined by maximum and minimum values expressed in binary with $q$ bits. This intuitively allows a broader range of discrete assignments for each input, while still operating within a discrete domain. (iii) The continuous setting, where $\mathcal{X}_{i}\subseteq\mathbb{R}$, which ensures guarantees across the entire infinite spectrum of the specified domain . Both discrete and continuous domains are widely used across various ML tasks, including those in vision and language tasks (Elkabetz and Cohen, 2021; Hong et al., 2022; Cartuyvels et al., 2021). We note that general discrete settings naturally encompass enumerable ones and are thus always at least as computationally challenging, whereas continuous settings do not enable a direct comparison to either and can make problems either harder or easier to solve (Karmarkar, 1984; Vavasis, 1991; Arora and Barak, 2009).
@@ 3.2 Dimension 2: The component model type
We consider a diverse range of component models that can be integrated into a GAM, spanning various levels of interpretability and reflecting commonly used approaches in the literature: (i) The classic spline-based models (Hastie, 2017), where each component function is piecewise polynomial. In practice, these polynomials are typically restricted to a degree of at most three, yielding the widely used cubic splines (Hastie, 2017), which will be our primary focus (though our results can often be extended to higher-order splines — see Appendix B). To distinguish these GAMs from the non-smooth variants, we refer to GAMs with spline-based components as Smooth GAMs. (ii) Neural Additive Models (NAMs)(Agarwal et al., 2021; Bouchiat et al., 2023; Xu et al., 2023; Jiao et al., 2024; Thielmann et al., 2024), in which each component is a neural network with ReLU activations. (iii) Explainable Boosting Machines (EBMs)(Caruana et al., 2015; Wei et al., 2022; Nori et al., 2021), where each component function is a boosted tree ensemble. We present the complete formal definitions of all model types in Appendix B.
@@ 3.3 Dimension 3: The explanation type
To explore various dimensions of GAM interpretability, we examine a broad range of widely used explanation types drawn from the existing literature. Building upon prior computational complexity frameworks for evaluating explanations in ML models (Barceló et al., 2020; Arenas et al., 2021a; Bassan et al., 2024), we formalize each type of explanation as an explainability query. Such a query takes a model $f$ and x as inputs and aims to address specific questions, offering a meaningful interpretation of the prediction $f(\textbf{x})$.
Sufficient reason Feature Selection. We consider the common sufficiency criterion for feature selection, used in popular explainability methods (Ribeiro et al., 2018; Carter et al., 2019; Ignatiev et al., 2019b). A sufficient reason is a subset of input features, $S\subseteq[n]$, such that fixing the features in $S$ to their values in $\textbf{x}\in\mathbb{F}$ guarantees the prediction remains $f(\textbf{x})$, regardless of the assignment to $\overline{S}$. We write $(\mathbf{x}_{S};\mathbf{z}_{\bar{S}})$ for an input where $S$ takes values from x and $\overline{S}$ from z. Formally, $S$ is a sufficient reason for $\langle f,\textbf{x}\rangle$ iff for all $\textbf{z}\in\mathbb{F}$: $f(\textbf{x}_{S};\textbf{z}_{\bar{S}})=f(\textbf{x})$.
A common assumption in the literature is that smaller sufficient reasons (those with smaller $|S|$) are more useful (Ribeiro et al., 2018; Carter et al., 2019; Ignatiev et al., 2019b). This motivates the search for cardinally minimal sufficient reasons, also called minimum sufficient reasons, and leads to our first explainability query:
@@
To deepen our understanding of the complexity of sufficient reasons, we also examine two common related explainability queries that refine the MSR query (Barceló et al., 2020; Bassan et al., 2024): (i) Check-Sufficient-Reason(CSR), which checks if a given subset $S$ is a sufficient reason; (ii) Count-Completions(CC), a generalized version of CSR that measures the fraction of completions preserving the prediction, capturing the probability of maintaining the classification. Full formalizations appear in Appendix C.
Contrastive Explanations. An alternative way to interpret models is by identifying subsets of features that, when changed, could alter the model’s prediction (Dhurandhar et al., 2018; Guidotti, 2022). We define s subset $S\subseteq[n]$ as contrastive if modifying its values can change the classification $f(\textbf{x})$, i.e., there exists $\textbf{z}\in\mathbb{F}$ such that $f(\textbf{x}_{\bar{S}};\textbf{z}_{S})\neq f(\textbf{x})$. As with sufficient reasons, smaller contrastive subsets are typically assumed to be more meaningful (Dhurandhar et al., 2018; Guidotti, 2022; Ignatiev et al., 2020; Mothilal et al., 2020), motivating a focus on cardinally-minimal contrastive reasons:
@@
Shapley Values. In the additive attribution setting, each feature $i\in[n]$ is assigned an importance weight $\phi_{i}$. A common method for assigning these weights is the Shapley value attribution index (Lundberg and Lee, 2017):
@@
We use the standard conditional expectation value function $v(S):=\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{z})\mid\textbf{z}_{S}=\textbf{x}_{S}]$ (Sundararajan and Najmi, 2020; Lundberg and Lee, 2017). Moreover, as is standard in both complexity analyses (Arenas et al., 2023; Van den Broeck et al., 2022) and practical SHAP methods like KernelSHAP (Lundberg and Lee, 2017), we assume feature independence. See Appendix C for a full formalization.
@@
Feature redundancy. For our final explanation form, we study the complexity of determining whether a feature $i$ is redundant with respect to a model $f$, following the notion of redundancy explored in prior work (Arenas et al., 2021a; Bassan et al., 2024; Huang et al., 2023; Darwiche and Hirth, 2020). Unlike the previous local explanation forms, this one is global — it checks if $i$ is redundant for all $\textbf{x}\in\mathbb{F}$. Formally, $i$ is redundant if for all $\textbf{x},\textbf{z}\in\mathbb{F}$, we have $f(\textbf{x})=f(\textbf{x}_{[n]\setminus\{i\}};\textbf{z}_{\{i\}})$.
@@ 4 Main Complexity Results
NAMs, EBMs coNP-H NP-C coNP-C #\#P-C #\#P-C
@@
NAMs, EBMs coNP-H NP-C coNP-C #\#P-C #\#P-C
@@ 5 A Computational Interpretability Hierarchy for GAMs
Our results uncover a rich and varied landscape of complexity outcomes across different settings, which we analyze in this section. To present these complexity distinctions more elegantly, we adopt the notation from (Barceló et al., 2020), indicating when one model is strictly more computationally interpretable (c-interpretable) than another — that is, when it belongs to a strictly lower complexity class.
@@ 5.2 The Significance of the Component Models on Complexity Depends on the Input Domain
We also note that the result shown in this subsection, together with our findings in Subsection 5.1 on input domain sensitivity, offers a complementary perspective to empirical work such as (Chang et al., 2021). These studies demonstrate that both the choice of component model and the form of data representation (i.e., input representation) substantially influence the learned functional structure of GAMs. Since differing functional forms naturally correspond to different levels of computational complexity, this aligns with our observation that the complexity of generating explanations is highly sensitive to these modeling choices.
@@ 5.3 Complexity Distinctions Between Regression and Classification, but Only for SHAP
We note that while SHAP allows for a direct comparison between regression and classification tasks, adapting our other explanation types — originally defined for classification — to regression requires redefining them (as in (Wu et al., 2023; Izza et al., 2024)) to assess whether values lie within a specified $\delta$-range. From a complexity standpoint, these behave similarly to classification cases and do not produce the same distinctions observed with SHAP. We elaborate on this in Appendix D.
@@
As noted in the introduction, the ML community often assumes that increasing a model’s additivity improves interpretability. For instance, NAMs (Agarwal et al., 2021; Radenovic et al., 2022) and EBMs (Caruana et al., 2015) are generally seen as more interpretable than standard neural networks and boosted trees, respectively. It thus seems natural to expect that explanations for additive models (e.g., NAMs) are easier to compute than their non-additive counterparts. Our findings support this, showing a clear complexity gap between black-box models and their additive counterparts. However, importantly, this gap is not universal — it strongly depends on the explanation type and the input domain. Table 2 summarizes prior complexity results for neural networks and tree ensembles, alongside our new results for NAMs and EBMs.
@@
MSR PTIME $\mathbf{\Sigma^{P}_{2}}$-C coNP-H, in $\Delta^{P}_{2}$ $\mathbf{\Sigma^{P}_{2}}$-C
@@ 6.1 Polynomial-Time Explanation Computations
Sufficient and contrastive reasons. We begin by analyzing the polynomial-time complexity of computing sufficient and contrastive reason queries (CSR, MSR, MCR). A key observation is that by identifying the minimal and maximal values each component $\beta_{i}\cdot f_{i}(\textbf{x}_{i})$ can take, we can rank features by their “importance”, measured by how much altering each feature deviates the prediction. A greedy algorithm, akin to those in (Barceló et al., 2020; Marques-Silva et al., 2020), then selects features by importance until a minimal sufficient or contrastive reason is found. CSR is simpler, requiring only a sufficiency check for a given subset. Thus, the main computational challenge lies in efficiently computing the min/max values of each $f_{i}(\textbf{x}_{i})$. For enumerable discrete inputs, this can be done by enumeration; for smooth GAMs, these values can be computed analytically, even with general discrete or continuous inputs.
@@
SHAP explanations in regression settings. When computing SHAP, polynomial-time computation is feasible only for GAMs in regression tasks (classification intractability is discussed in Subsection 6.3). Leveraging the linearity axiom of Shapley values (Lundberg and Lee, 2017), we reduce the computation to evaluating the computations of $\beta_{i}\cdot\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f_{i}(\textbf{x}_{i})]$. When the model is either discretely enumerable or supports efficient integration (e.g., in Smooth GAMs), these expectations can be computed in polynomial time.